DEPERTURBATION STUDIES OF AlO : INTERACTIONS IN THE A^{2}Π ∼ X^{2}Σ^{+} STATES

\underline{K. SUNANDA}, M. D.SAKSENA, and B. N. JAGATAP, Atomic and Molecular Physics Division, Bhabha Atomic Research Centre, Trombay, Mumbai 400 085,India; M. N. DEO, High Pressure Physics Division, Bhabha Atomic Research Centre, Trombay, Mumbai 400 085, India; N. MHASKE and S. H. BEHERE, Department of Physics, Dr. Babasaheb Ambedkar Marathwada University, Aurangabad, India.

The rotational structure of D^{2}Σ^{+} - A^{2}Π band system of AlO molecule, at moderate high resolution, was photographed by Singh et. al^{a}. They carried out the rotational analyses of seven bands of this system, involving vibrational levels v′= 0 - 5 and v′′= 0 - 4 for both the sub-transitions D^{2}Σ^{+} - A^{2}Π_{1/2} and D^{2}Σ^{+} - A^{2}Π_{3/2} and reported several rotational perturbations in the A^{2}Π_{i} state. In this work, we present a deperturbation analysis which yields deperturbed molecular constants of the X^{2}Σ^{+}, A^{2}Π_{i}, and D^{2}Σ^{+} states. The revised molecular parameters for the X^{2}Σ^{+} and D^{2}Σ^{+} states are first obtained from a simultaneous fit performed using previous high resolution data of the D^{2}Σ^{+} - X^{2}Σ^{+} system ^{b} involving v′=0 - 6 and v′′= 0 - 4 vibrational levels and B^{2}Σ^{+} - X^{2}Σ^{+}, system ^{c} involving v′= 0 - 11 and v′′= 0 - 7 vibrational levels. PGOPHER program is used to simulate and fit the observed spectra^{d} . The branch frequencies involving the D^{2}Σ^{+} - A^{2}Π_{i} transitions from our earlier studies have been incorporated into a single Hamiltonian to obtain improved molecular constants together with the L- doubling and spin splitting coefficients for these states. Further, invoking perturbing state (X^{2}Σ^{+}) molecular parameters in this fit, deperturbation of the vibrational levels within the A^{2}Π_{i} state of AlO up to v′′ = 4 is obtained. A global least squares fit to all the data allows determination of A^{2}Π_{i} state molecular constants with much improved precision. The results of this study will be presented.

aM. Singh and M. D. Saksena, Can. J. Phys. 63, 1162 (1985).

bM. Singh and M. D. Saksena, Proc. Indian Acad. Sci. 77, 139 (1973).

cM. D. Saksena, M. N. Deo, K. Sunanda, S. H. Behere and C. T. Londhe, 247, 47(2008).

dPGOPHER, a Program for Simulating Rotational Structure, C. M.Western, University of Bristol.